OpenStudy (anonymous):

x − 7y = 20 x + 12y = −1 Find the value for x

OpenStudy (2bornot2b):

-19y=21 => y=-21/19 Now you can solve x

OpenStudy (anonymous):

subtract columns 0x - 19y = 21 or -19y = 21 which means y = 21/-19 substitute into x + 12y = −1 x + 12(21/-19) = −1 or x - 252/21 = -1 add 252/21 to both sides therefore, x = 231/21